Given f(x)=x2-4x-5 defined for x>2. Find (f-1)'(0). Also, we know f(5)=0
Answer #1

Given function, f(x)=x2 - 4x-5 Let f (x)=y=x2 - 4x-5 x2 - 4x = y +5 x2 - 4x +4=y+5+4 (x-2)* = y+9 x-2=+Xy+9 x=2+y+9 f(y)=2+/y

1 1 Our object is to find the derivative using power rule of the given function: So derivative with respect to x on both side